# Tree Ensembles for Contextual Bandits

### v1

@@ Abstract
We propose a novel framework for contextual multi-armed bandits based on tree ensembles. Our framework integrates two widely used bandit methods, Upper Confidence Bound and Thompson Sampling, for both standard and combinatorial settings. We demonstrate the effectiveness of our framework via several experimental studies, employing XGBoost, a popular tree ensemble method. Compared to state-of-the-art methods based on neural networks, our methods exhibit superior performance in terms of both regret minimization and computational runtime, when applied to benchmark datasets and the real-world application of navigation over road networks.
@@ 1 Introduction
Stochastic multi-armed bandits (MAB) (see e.g., Slivkins, 2019) provide a principled framework for making optimal sequences of decisions under uncertainty. An important variant known as the contextual multi-armed bandit incorporates additional contextual / side information into the decision-making process, allowing for more personalized or adaptive action selection. Following the recent success of (deep) neural networks in solving various machine learning tasks, several methods building on such models have been suggested for finding functional relationships between the context and outcomes of actions to aid the decision-making process (e.g., Zhou et al., 2020; Zhang et al., 2021; Zhu and Van Roy, 2023; Osband et al., 2021). Although delicate, these methods can be impractical to implement, time-consuming, or expensive to use.
In light of these challenges, we propose bandit algorithms that utilize tree ensemble (TE) methods (see Hastie et al., 2009), such as gradient-boosted decision trees (GBDT) and Random Forests, to comprehend the contextual features, combined with the most popular bandit methods Upper Confidence Bound (UCB) and Thompson Sampling (TS). Compared with neural network methods, our methods, called TEUCB and TETS, yield superior results on UCI benchmark datasets. Furthermore, our methods benefit from more effective learning and less computational overhead. We also extend our methods to the combinatorial contextual setting, a variant of bandits that deals with complex combinatorial sets of arms, and investigate them on the important real-world application of navigation through road networks.
@@ 1.1 Related Work
The concept of applying decision trees to contextual bandit problems has only been studied in a limited way. Elmachtoub et al. (2017) suggest using a single decision tree for each arm available to the agent. One issue with this approach is that it might have trouble reaching the accuracy and robustness of tree ensemble methods, especially when the objective is to estimate complex reward functions. Furthermore, storing and fitting one model per arm does not scale well with a large action space (e.g., in the combinatorial setting), and may also lead to excessive exploration for dynamical arm sets, as it cannot attend to what it has learned from the contexts of other arms. In contrast, Féraud et al. (2016) employs an ensemble of trees, where they use decision trees of unit depth, also known as decision stumps. While tree ensembles with a large number of shallow trees tend to perform well regarding both accuracy and robustness, limiting the depth in such a way (i.e., too extremely) is often not an optimal choice. Additionally, Féraud et al. only investigate the bandit algorithm known as successive elimination, which usually leads to less efficient exploration than UCB and TS, and thus is less commonly used. Moreover, neither Elmachtoub et al. (2017) nor Féraud et al. (2016) consider combinatorial contextual bandits.
@@
Apart from the two works mentioned above, the problem has primarily been addressed using other machine learning models. When the reward function is linear w.r.t. the context, LinUCB (Li et al., 2010) and Linear Thompson Sampling (LinearTS) (Agrawal and Goyal, 2013) have shown good performance. In more complicated cases, the linear assumption does not hold anymore, and thus more expressive models are needed. In recent years, neural contextual bandits have attracted a lot of attention. As the name suggests, these methods use (deep) neural networks to model the expected rewards for some set of observed contextual features. Out of several proposed neural contextual bandits, NeuralUCB (Zhou et al., 2020) and NeuralTS (Zhang et al., 2021), in particular, have been shown to perform well on regret minimization tasks, while also providing theoretical guarantees through regret bounds. On the negative side, due to the way these methods estimate the uncertainty in predictions, they rely on inverting a matrix of size equal to the number of network parameters, which is usually computationally inefficient and time-consuming.
@@
Zhu and Van Roy (2023) provide a thorough review of different neural bandit methods, and suggest a method based on the work by Osband et al. (2021) for more efficient uncertainty modeling, showing its advantages in terms of regret minimization and computational efficiency. Despite their promising results, we argue that neural networks may not necessarily be the most sophisticated models for contextual bandits. For instance, even with epistemic neural networks, training is still resource-intensive. On the other hand, there is a large body of work demonstrating the effectiveness of tree ensemble methods for both regression and classification tasks in the standard supervised learning paradigm (Borisov et al., 2022; Gorishniy et al., 2021; Grinsztajn et al., 2022), and thereby, we believe that their extension to contextual bandits can be of great potential.
@@
One less directly related method is Ensemble Sampling (Lu and Van Roy, 2017), where the ensembles are solely used for uncertainty estimation and not for modeling the reward functions in bandits. Uncertainty estimates in Ensemble Sampling and similar methods focus on uncertainties in the parameters of the models used to approximate the underlying processes being sampled. In contrast, we sample directly from estimated distributions of the expected rewards given contexts. Our approach aligns closer with sampling-based bandit methods like LinearTS and NeuralTS.
@@ 2.1 Multi-Armed Bandit Problem
The multi-armed bandit (MAB) problem is a sequential decision-making problem under uncertainty, where the decision maker, i.e., the agent, interactively learns from the environment over a horizon of $T$ time steps (interactions). At each time step $t\leq T$, the agent is presented with a set of $K$ actions $\mathcal{A}$, also referred to as arms. Each action $a\in\mathcal{A}$ has an (initially) unknown reward distribution with expected value $\mu_{a}$. The objective is to maximize the cumulative reward over the time horizon $T$, or more commonly (and equivalently) to minimize the cumulative regret.,which is computed as
@@ 2.2 Contextual Bandit Problem
The contextual multi-armed bandit problem is an extension of the classical MAB problem described in Section 2.1. The agent is, at each time step $t$, presented with a context vector $\mathbf{x}_{t,a}\in\mathbb{R}^{d}$ for each action $a\in\mathcal{A}$. For example, a recommendation system could encode a combination of user-related and action-specific features into the context vector $\mathbf{x}_{a}$. Then, the expected reward of action $a$ at time $t$ is given by an (unknown) function of the context $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, such that $\mathbb{E}[r_{t,a}]=f(\mathbf{x}_{t,a})$. Learning to generalize the relationship between the contextual features and the expected reward is crucial for effective learning and minimizing the regret.
@@ 3 Proposed Algorithms
Despite the potential benefits, these types of models have not been studied much for contextual bandits, and to the best of our knowledge, there are no previously known works that combine tree ensemble models in a principled way with the most effective methods known for handling the exploration-exploitation dilemma. We address this research gap and introduce a novel approach for the contextual MAB problem that combines any tree ensemble method with a natural adaption of the two most prominent exploration schemes, UCB and TS. Within our framework, we empirically investigate the XGBoost algorithm and demonstrate its very promising performance on a number of benchmark tasks, as well as a combinatorial contextual MAB problem for efficient navigation over a stochastic road network in Section 4. However, we emphasize that our bandit algorithms are sufficiently general to be employed together with any type of decision tree ensemble.
@@ 3.1 Weak Learners
where the outputs $o_{i}$, in turn, are averages of the output suggested by each training sample based on its target value.
@@ 3.2 Uncertainty Modeling
We acknowledge that these assumptions may not be always true in practice, but they act here as motivation for the design of our proposed algorithms. Specifically, they entail a straightforward way of accumulating the variances of the individual tree contributions to obtain an uncertainty estimate in the total reward prediction, allowing us to construct efficient exploration strategies. In the following two subsections, we present our approach to doing so with UCB and TS methods, respectively.
@@ 3.3 Tree Ensemble Upper Confidence Bound
UCB methods act under the principle of optimism in the face of uncertainty, and have established themselves as one of the most prominent approaches to handling exploration in bandit problems. A classic example is the UCB1 algorithm (Auer et al., 2002) which builds confidence bounds around the expected reward from each arm depending on the fraction of times they have been played, and for which there are proven upper bounds on the expected regret. One disadvantage of the method, however, is that it does not take the variance of the observed rewards into account, which may lead to a sub-optimal exploration strategy in practice. In light of this, Auer et al. (2002) further proposed the UCB1-Tuned and UCB1-Normal algorithms, which extend UCB1 by including variance estimates, and demonstrated better performance experimentally. The main difference between the two extended versions is that UCB1-Tuned assumes Bernoulli-distributed rewards, while UCB1-Normal is constructed for Gaussian rewards. At each time step $t$, UCB1-Normal selects the arm $a$ with the maximal upper confidence bound $U_{t,a}$, calculated as
@@
By the assumptions we have made on the tree ensembles, the predictions of the expected rewards will be approximately Gaussian. Therefore, we propose an algorithm called Tree Ensemble Upper Confidence Bound (TEUCB) in Algorithm 1, which draws inspiration from UCB1-Normal, and is suitable for both Gaussian and Bernoulli bandits. As seen in lines 21 and 23 of Algorithm 1, the selection rule of TEUCB closely resembles that of UCB1-Normal, but is constructed specifically for contextual MABs where the arms available in each time step are characterized by their context vectors. Therefore, TEUCB considers each individual contribution from all data points in a leaf as a sample of that leaf’s output distribution, which is demonstrated in line 13. The total prediction for a context $\mathbf{x}_{t,a}$ is subsequently computed as the sum of sampled contributions from each leaf that $\mathbf{x}_{t,a}$ is assigned to in the ensemble (lines 15, 16, 17).
@@
21:        $a_{t}\leftarrow\text{argmax}_{a}U_{t,a}$
22:     end for
@@ 3.4 Tree Ensemble Thompson Sampling
19:        $a_{t}\leftarrow\text{argmax}_{a}\tilde{r}_{t,a}$
20:     end for
@@ 3.5 Extension to Combinatorial Bandits
Algorithm 1 and Algorithm 2 are formulated to address the standard contextual MAB problem. However, TEUCB and TETS can easily be extended to the combinatorial semi-bandit setting. The main difference is in the way arms are selected, i.e., as super arms instead of individual arms. In Algorithm 1, this corresponds to modifying line 21 to
@@
The selected super arm $\mathcal{S}_{t}$ is subsequently played in line 23. Algorithm 2 can be similarly modified in lines 19 and 21, respectively. Furthermore, the set of observed context-reward pairs would include all rewards received for each base arm in the selected super arms individually, which are generally more than one per time step.
@@ 3.6 XGBoost Adaptation
By using XGBoost regression models, we can extract $p$ directly from each individual leaf in the constructed ensemble. The sample variance $s^{2}$ is not accessible directly, but we can easily calculate it. As a GBDT method, XGBoost calculates its outputs as the average difference between the target value and the value predicted from the collection of preceding trees in the ensemble, multiplied with a learning rate. Hence, the trees do not have predetermined weights as in, e.g., Random Forests. Instead, the relative size of the contributions of an assigned leaf to the final prediction depends on the particular paths that are traversed through the other trees, which may be different for every sample we observe in a leaf. Therefore, we cannot estimate $s^{2}$ from the variance in target values directly.
However, XGBoost is augmented with many useful features, one of them being staged predictions (XGBoost Developers, 2022a). This means that we can propagate the predicted outputs on sub-ensembles up to a certain tree on all data samples and cache these outputs. After recording these we include the next tree in the prediction as well, without having to start over. By utilizing this technique, $s^{2}$ is easily estimated from the samples assigned to a particular leaf. Furthermore, $c$ is simply the number of such samples. This gives us all the pieces of the puzzle needed to apply XGBoost to the TEUCB and TETS algorithms.

### v2

@@ Abstract
We propose a novel framework for contextual multi-armed bandits based on tree ensembles. Our framework integrates two widely used bandit methods, Upper Confidence Bound and Thompson Sampling, for both standard and combinatorial settings. We demonstrate the effectiveness of our framework via several experimental studies, employing both XGBoost and random forest, two popular tree ensemble methods. Compared to state-of-the-art methods based on decision trees and neural networks, our methods exhibit superior performance in terms of both regret minimization and computational runtime, when applied to benchmark datasets and the real-world application of navigation over road networks.
@@ 1 Introduction
Stochastic multi-armed bandits (MABs) (Slivkins, 2019) provide a principled framework for making optimal sequences of decisions under uncertainty. As discussed by Zhu and Van Roy (2023), MABs constitute a vital component of modern recommendation systems for efficient exploration, among many other applications. An important variant known as the contextual multi-armed bandit incorporates additional contextual / side information into the decision-making process, allowing for more personalized or adaptive action selection. Following the recent success of (deep) neural networks in solving various machine learning tasks, several methods building on such models have been suggested for finding functional relationships between the context and outcomes of actions to aid the decision-making process (Zhou et al., 2020; Zhang et al., 2021; Zhu and Van Roy, 2023; Osband et al., 2021). Although delicate, these methods can be impractical, time-consuming, and computationally expensive.
In light of these challenges, we propose bandit algorithms that utilize tree ensemble (TE) methods (Hastie et al., 2009), like gradient-boosted decision trees (GBDT) and random forests, to comprehend the contextual features, combined with the most popular bandit methods Upper Confidence Bound (UCB) and Thompson Sampling (TS) as exploration strategies. Compared with both other tree-based and neural network methods, our methods, called TEUCB and TETS, yield superior results on UCI benchmark datasets. Furthermore, our methods benefit from more effective learning with less computational overhead. We furthermore extend the framework to the combinatorial contextual setting, a variant of bandits that deals with complex combinatorial sets of arms, and investigate them on the important real-world application of efficient navigation in road networks.
@@ 1.1 Related Work
The concept of applying decision trees to contextual bandit problems has been studied to a limited extent. Elmachtoub et al. (2017) suggest using a single decision tree for each arm available to the agent. Their method is referred to as TreeBootstrap, as the concept of bootstrapping is employed in order to resemble the Thompson Sampling process. One issue with their method is that it might have trouble reaching the accuracy and robustness of tree ensemble methods, especially when estimating complex reward functions. The possibility of applying random forests in their framework is discussed, but no concrete methods or experimental results are presented. Furthermore, storing and training one model per arm does not scale, especially not with large action spaces (e.g., in the combinatorial setting), and could potentially lead to excessive exploration for dynamic arm sets, as it cannot attend to what it has learned from the contexts of other arms.
In contrast, Féraud et al. (2016) employ a tree ensemble but only uses decision trees of unit depth also known as decision stumps. While tree ensembles with a large number of decision stumps tend to perform well regarding accuracy, limited variance, and increased robustness, restricting the depth in such an extreme way may not be adequate when addressing complex tasks (Hastie et al., 2009). The experimental studies by Hastie et al. (2009) indicate that tree depths of 4 to 8 work well for boosting. Additionally, Féraud et al. (2016) only investigate the bandit algorithm known as successive elimination, which usually leads to less efficient exploration compared to UCB and TS, and thus is used less commonly. Moreover, the method, called Bandit Forest, requires binary encoded contextual features. Comparing these two tree-based bandit methods, Elmachtoub et al. (2017) experimentally show that TreeBootstrap tends to outperform Bandit Forest in practice.
It is worth noting that neither Elmachtoub et al. (2017) nor Féraud et al. (2016) consider combinatorial contextual bandits. Apart from these two methods, the problem has primarily been addressed using other machine learning models. When the reward function is linear w.r.t. the context, LinUCB (Li et al., 2010) and Linear Thompson Sampling (LinTS) (Agrawal and Goyal, 2013) have demonstrated good performance. In more complicated cases, the linear assumption no longer holds, and thus more expressive models are needed.
@@
In recent years, neural contextual bandits have attracted a lot of attention. As the name suggests, these methods use (deep) neural networks to model the expected rewards for some set of observed contextual features. Out of several proposed neural contextual bandit algorithms, NeuralUCB (Zhou et al., 2020) and NeuralTS (Zhang et al., 2021), in particular, have been shown to perform well, while also providing theoretical regret bounds. On the negative side, due to how these methods estimate the uncertainty in predictions, they rely on inverting a matrix of size equal to the number of network parameters, which is usually computationally expensive and time-consuming, especially when the prediction tasks require large neural networks.
@@
Zhu and Van Roy (2023) provide a thorough review of different neural bandit methods and suggests a method based on epistemic neural networks (Osband et al., 2021) for more efficient uncertainty modeling, showing its advantages in terms of regret minimization and computational efficiency. Despite their promising results, we argue that neural networks may not necessarily be the most appropriate models for contextual bandits. For instance, even with epistemic neural networks, training is still resource-intensive. On the other hand, there is a large body of work demonstrating the effectiveness of tree ensemble methods for both regression and classification tasks in the standard supervised learning paradigm (Borisov et al., 2022; Gorishniy et al., 2021; Grinsztajn et al., 2022), and thereby, we believe that their extension to contextual bandits can be of great potential.
@@
One less directly related method is Ensemble Sampling (Lu and Van Roy, 2017), where the ensembles are solely used for uncertainty estimation and not for modeling the reward functions in bandits. Uncertainty estimates in Ensemble Sampling and similar methods focus on uncertainties in the parameters of the models used to approximate the underlying processes being sampled. In contrast, we sample directly from estimated distributions of the expected rewards given contexts. Our approach aligns closer with sampling-based bandit methods like LinTS and NeuralTS.
@@ 2.1 Multi-Armed Bandit Problem
The multi-armed bandit (MAB) problem is a sequential decision-making problem under uncertainty, where the decision maker, i.e., the agent, interactively learns from the environment over a horizon of $T$ time steps (interactions). At each time step $t\leq T$, the agent is presented with a set of $K$ actions $\mathcal{A}$, commonly referred to as arms. Each action $a\in\mathcal{A}$ has an (initially) unknown reward distribution with expected value $\mu_{a}$. The objective is to maximize the cumulative reward over the time horizon $T$, or more commonly (and equivalently) to minimize the cumulative regret, which is calculated as
@@ 2.2 Contextual Bandit Problem
The contextual multi-armed bandit problem is an extension of the classical MAB problem described in Section 2.1. The agent is, at each time step $t$, presented with a context vector $\mathbf{x}_{t,a}\in\mathbb{R}^{d}$ for each action $a\in\mathcal{A}$. For example, a recommendation system could encode a combination of user-related and action-specific features into the context vector $\mathbf{x}_{a}$. Then, the expected reward of action $a$ at time $t$ is given by an (unknown) function of the context $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, such that $\mathbb{E}[r_{t,a}]=f(\mathbf{x}_{t,a})$. Learning to generalize the relationship between the contextual features and the expected reward is crucial for effective learning and minimizing the cumulative regret.
@@ 3 Proposed Algorithms
Despite the potential benefits, these types of models have not been studied much for contextual bandits. To the best of our knowledge, there are no previously known works that in a principled way combine tree ensemble models with UCB, which is one of the most effective strategies known for handling the exploration-exploitation dilemma. Further, works that combine tree ensemble methods with Thompson Sampling are limited. We address this research gap and introduce a novel approach for the contextual MAB problem that combines any tree ensemble method with a natural adaption of these two prominent exploration schemes.
Within our framework, we empirically investigate the XGBoost and random forest algorithms and demonstrate their auspicious performance on several benchmark tasks, as well as a combinatorial contextual MAB problem for efficient navigation over a stochastic road network in Section 4. However, we emphasize that our bandit algorithms are sufficiently general to be employed with any decision tree ensemble.
@@ 3.1 Weak Learners
where the outputs $o_{i}$, in turn, are averages of the output suggested by each training sample based on its target value. We use this definition of $p$ to unify the notation of all kinds of tree ensembles. Hence, for averaging models, such as random forest, we assume that the leaf values $o_{i}$ are already weighted by $1/N$.
@@ 3.2 Uncertainty Modeling
We acknowledge that these assumptions may not always be true in practice, but they act here as motivation for the design of our proposed algorithms. Specifically, they entail a straightforward way of accumulating the variances of the individual tree contributions to obtain an uncertainty estimate in the total reward prediction, allowing us to construct efficient exploration strategies. In the following two subsections, we present our approach to doing so with UCB and TS methods, respectively.
@@ 3.3 Tree Ensemble Upper Confidence Bound
UCB methods act under the principle of optimism in the face of uncertainty, and have established themselves as one of the most prominent approaches to handling exploration in bandit problems. A classic example is the UCB1 algorithm (Auer et al., 2002) which builds confidence bounds around the expected reward from each arm depending on the fraction of times they have been played, and for which there are proven upper bounds on the expected regret. One disadvantage of the method, however, is that it does not take the variance of the observed rewards into account, which may lead to a sub-optimal exploration strategy in practice. In light of this, Auer et al. (2002) further proposed the UCB1-Tuned and UCB1-Normal algorithms, which extend UCB1 by including variance estimates, and demonstrate better performance experimentally. The main difference between the two extended versions is that UCB1-Tuned assumes Bernoulli-distributed rewards, while UCB1-Normal is constructed for Gaussian rewards. At each time step $t$, UCB1-Normal selects the arm $a$ with the maximal upper confidence bound $U_{t,a}$, calculated as
@@
By the assumptions we have made on the tree ensembles, the predictions of the expected rewards will be approximately Gaussian. Therefore, we propose an algorithm called Tree Ensemble Upper Confidence Bound (TEUCB) in Algorithm 1, which draws inspiration from UCB1-Normal, and is suitable for both Gaussian and Bernoulli bandits. As seen in lines 22 and 23 of Algorithm 1, the selection rule of TEUCB closely resembles that of UCB1-Normal, but is constructed specifically for contextual MABs where the arms available in each time step are characterized by their context vectors. Therefore, TEUCB considers each individual contribution from all data points in a leaf as a sample of that leaf’s output distribution, which is demonstrated in line 13. The total prediction for a context $\mathbf{x}_{t,a}$ is subsequently computed as the sum of sampled contributions from each leaf that $\mathbf{x}_{t,a}$ is assigned to in the ensemble (lines 15, 16, 17).
@@
21:     end for
22:     $a_{t}\leftarrow\text{argmax}_{a}U_{t,a}$
@@ 3.4 Tree Ensemble Thompson Sampling
19:     end for
20:     $a_{t}\leftarrow\text{argmax}_{a}\tilde{r}_{t,a}$
@@ 3.5 Extension to Combinatorial Bandits
Algorithm 1 and algorithm 2 are formulated to address the standard contextual MAB problem. However, TEUCB and TETS can easily be extended to the combinatorial semi-bandit setting. The main difference is in the way arms are selected, i.e., as super arms instead of individual arms. In Algorithm 1, this corresponds to modifying line 22 to
@@
The selected super arm $\mathcal{S}_{t}$ is subsequently played in line 23. Algorithm 2 can be similarly modified in lines 20 and 21, respectively. Furthermore, the set of observed context-reward pairs would include all rewards received for each base arm in the selected super arms individually, which are generally more than one per time step.
@@ 3.6 XGBoost Adaptation
By using XGBoost regression models, we can extract $o$ directly from each individual leaf in the constructed ensemble. The sample variance $s^{2}$ is not accessible directly, but we can easily calculate it. As a GBDT method, XGBoost calculates its outputs as the average difference between the target value and the value predicted from the collection of preceding trees in the ensemble, multiplied with a learning rate. Hence, the trees do not have predetermined weights as in, e.g., random forests. Instead, the relative size of the contributions of an assigned leaf to the final prediction depends on the particular paths that are traversed through the other trees, which may be different for every sample we observe in a leaf. Therefore, we cannot estimate $s^{2}$ from the variance in target values directly.
However, XGBoost is augmented with many useful features, one of them being staged predictions (XGBoost Developers, 2022a). This means that we can propagate the predicted outputs on sub-ensembles up to a certain tree on all data samples and cache these outputs. After recording these we include the next tree in the prediction as well, without having to start over. By utilizing this technique, $s^{2}$ is easily estimated from the previously observed samples assigned to a particular leaf. Furthermore, $c$ is simply the number of such samples. This gives us all the pieces of the puzzle needed to apply XGBoost to the TEUCB and TETS algorithms. In order to reduce bias when calculating $o$, $s^{2}$ and $c$, one may split the previously observed samples into two distinct data sets; one which is used for building the trees, and a second for the value estimations.
3.7 Random Forest Adaptation
Similar to XGBoost, random forests can also be incorporated into TEUCB and TETS. This is even less complicated in the case of random forests as the trees in a random forest do not depend on the predictions made by any of the other trees. Therefore, $s^{2}$ can be directly estimated from the variance in the target values, taking the tree’s weight of $1/N$ into account. Hence, all the quantities of interest (i.e., $o$, $s^{2}$, and $c$) are available from the trees of a random forest by looking at which observed data points are assigned to the terminal nodes.
As with XGBoost, one may consider dividing the previously observed samples into different data sets for fitting and value estimation. Alternatively, another way of handling bias when using random forests is to consider only the out-of-bag samples for computing $o$, $s^{2}$, and $c$. The random forest algorithm employs the concept of bagging—resampling with replacement—when building trees, which means roughly one third of the samples will be omitted in the construction of any particular tree. Therefore, random forests yield a natural way for providing an independent set of data points for model evaluation without limiting the size of the training data set.